OpenStudy (anonymous):

I NEED A QUICK ANSWER PLEASE!! Find the perimeter of an object which has a length of square root of 180 feet and a width of square root of 125 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. What type of object might this be? PLZ HELP MEE

OpenStudy (shamim):

Wt do u know abt perimeter

OpenStudy (error1603):

√180 = √(36*5) = √36 * √5 = 6√5 √125 = √(25*5) = √25 * √5 = 5√5 perimeter of a rectangular object = 2*√180 + 2*√125 = 2*6√5 + 2*5√5 = 12√5 + 10√5 = 22√5.

OpenStudy (anonymous):

So lets try few guess what could be the shape. Circle nope- it has only radius Rhombus- buts it's all side are equal and here it is given two side. Square - same as rhombus Rectange- It should have length and breadth. It fits the question. Perimeter of rectangle = 2*(length + width). Now find it?

OpenStudy (shamim):

Perimeter is a total length of the rectangle

OpenStudy (error1603):

final solution in simplified radical form = 22√5

OpenStudy (shamim):

U hv 2 length nd 2 weidth of a rectangle

OpenStudy (anonymous):

ty

OpenStudy (error1603):

√180 is about 13.4 feet √125 is about 11.2 feet

OpenStudy (anonymous):

ty very much